# Secure numerical simulations using fully homomorphic encryption

### v2

@@ 2.1.4 Addition
This operation results in an error growth from $e$ to $e+\hat{e}$ in decrypted data, which is still small compared to the data itself. To add a plaintext to a ciphertext, the plaintext is encoded and then extended to the ciphertext space (no additional error is introduced by this operation). Addition of ciphertexts and scalars is done similarly (but no encoding is needed in this case).
@@ 2.1.5 Key Switching
Unlike addition, the two other basic arithmetic operations, multiplication and rotation, require executing a special maintenance procedure called key switching. This key switching procedure needs an additional public key, which is often called an evaluation key or key-switching hint/key. We will use the term evaluation key in the paper.
@@
Note that key switching is more computationally expensive in CKKS than the actual multiplication or rotation, and is typically the main bottleneck in FHE applications based on CKKS.
@@ 2.1.6 Multiplication
Multiplication of two CKKS ciphertexts $\bm{c}$ and $\hat{\bm{c}}$ is a considerably more complex operation than addition, which is why we will only provide a brief overview here. It requires three steps:
@@
As we saw in Sec. 2.1.3, after encryption the first polynomial of the ciphertext $c_{0}$ is linear with respect to the secret key $\bm{\mathrm{sk}}$. After multiplying two ciphertexts using the tensor product, we get a quadratic polynomial in secret key polynomial $s$, i.e., the decryption would be evaluated as $\langle\bm{d},(1,s,s^{2})\rangle=d_{0}+d_{1}\cdot s+d_{2}\cdot s^{2}$. Each subsequent multiplication will increase the degree of $s$ even further; one can think of a multiplication of two ciphertexts as a product of two corresponding decryption polynomials. This implies the size of the ciphertext after a multiplication of two ciphertexts of sizes $i$ and $j$ will grow to $i+j-1$ polynomials. To compress the ciphertext from three to two polynomials (or decrement the size by one in a more general case), the relinearization procedure is needed. In summary, the main goal of relinearization is to maintain the compact representation of ciphertexts.
@@
In leveled FHE, one therefore tries to minimize the required multiplicative depth of an algorithm by evaluating a chain of multiplications using a binary tree approach. For example, instead of computing $\bm{c^{\prime}}=\bm{c}_{1}\cdot\bm{c}_{2}\cdot\bm{c}_{3}\cdot\bm{c}_{4}$ with three consecutive multiplications (and thus requiring three levels), one can first compute $\bm{r}_{1}=\bm{c}_{1}\cdot\bm{c}_{2}$ and $\bm{r}_{2}=\bm{c}_{3}\cdot\bm{c}_{4}$ using one level, and then combine them to $\bm{c^{\prime}}=\bm{r}_{1}\cdot\bm{r}_{2}$ using a second level. In the naïve case, to multiply $n$ ciphertexts, the multiplicative depth of $n-1$ is needed. If the binary tree method is used instead, the multiplicative depth is reduced to approximately $\log_{2}(n)$ (see also [38]).
@@ 2.1.7 Rotation
The multiplication and addition operations are element-wise, i.e., they are always performed between corresponding data elements in the user-provided data vectors. These operations are evaluated in a Single-Instruction-Multiple-Data (SIMD) manner, where one homomorphic operation computes many thousands of scalar additions/multiplications at once. Often, we need to perform operations over elements of the same vector or access data elements at particular indices. In these cases, the rotation operation is required, which cyclically shifts encrypted data by some rotation index. When the rotation is evaluated, the underlying secret key gets also rotated. To change the underlying secret key back to the original one (to support further homomorphic computations), one needs to apply the key switching operation. Note that the rotation itself is cheap (effectively it is just reindexing), but the required key switching operation following it is computationally expensive.
@@ 2.1.8 Bootstrapping
As described in Sec. 2.1.2 and 2.1.6, the necessity to rescale the ciphertext limits the number of multiplicative operations that can be performed, since no more multiplications are possible once the level $l$ reaches zero. The bootstrapping operation alleviates this restriction by resetting the level to $l>0$. It involves approximating a homomorphic decryption procedure on a ciphertext using an encrypted version of the secret key (this is done indirectly in CKKS), and then re-encrypting it to obtain a refreshed ciphertext. In theory, the number of multiplications that can be performed with a ciphertext is unlimited when using bootstrapping. However, bootstrapping also adds a significant approximation error to the ciphertext and is computationally very expensive. Due to the high complexity of the bootstrapping operation, we will not go into more detail here. In-depth descriptions of various bootstrapping implementations are covered elsewhere [41].
@@ 2.2 Software libraries for fully homomorphic encryption
As we have seen, the CKKS scheme is a powerful method for FHE with real numbers. At the same time, it is also an advanced technique that requires a deep understanding of the underlying mathematical methods to be used effectively. It is therefore advisable to use a software library that provides an implementation of the CKKS scheme and other necessary operations. Several such libraries are available, e.g., SEAL [42], HElib [43], and OpenFHE [40].
In this paper, we will focus on the OpenFHE library, which is actively maintained and includes most recent CKKS optimizations [44]. We will also present two Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface to OpenFHE in the Julia programming language. We will now give a brief overview of these tools.
@@ 2.2.1 OpenFHE
OpenFHE [40] is an open-source FHE library that offers efficient C++ implementations of all common FHE schemes, including the CKKS scheme with bootstrapping. OpenFHE contains many examples that are especially useful for new users when designing their applications. The OpenFHE library puts an emphasis on usability. For example, the relinearization and rescaling operations after ciphertext multiplication are handled automatically, allowing the library to be used also by non-experts in FHE. This can also be observed in a short code snippet in Listing 2.2.3 (left), where the CKKS internals for multiplication are hidden from the user. Note that some code for initializing the setup was omitted for clarity, e.g., to create the cryptographic context object (named cc in this listing) or the public and private keys. For more details about OpenFHE, please refer to [40].
@@
It is this feature that we leverage in our OpenFHE.jl package [46], which provides a Julia wrapper for the OpenFHE library. When installing OpenFHE.jl, pre-built binaries of the OpenFHE library are automatically downloaded without the user having to compile anything locally. The C++ functionality is exposed in Julia via the CxxWrap.jl package [47]. Besides offering a native Julia interface to the OpenFHE library, OpenFHE.jl does not provide any extra functionality. It is intended to be used as a backend for other Julia packages that require FHE capabilities. As can be seen in Listing 2.2.3 (right), the syntax of OpenFHE.jl is very similar to that of OpenFHE, making it easy to port code between the two programming languages.
@@
To make the use of FHE more accessible to non-experts in Julia, we created the SecureArithmetic.jl package [48]. It inherits the functionalities of the OpenFHE.jl package, but provides them consistently through a more convenient Julia syntax. Using the SecureArithmetic.jl package, users can write arithmetic expressions in a common mathematical notation, significantly simplifying the code. Furthermore, SecureArithmetic.jl supports the ability to use the same code for secure and non-secure computations, which is especially useful during the design stage. It allows one to debug FHE algorithms without having to go through the actual encryption and decryption steps, making the execution much faster during the prototyping phase. Listing 2.2.3 shows an example of how the ciphertext multiplication is performed using SecureArithmetic.jl. Compared to OpenFHE/OpenFHE.jl in Listing 2.2.3, SecureArithmetic.jl provides a more high-level interface that is closer to the mathematical notation. However, since it is built on top of OpenFHE.jl and OpenFHE, it is still possible to access the lower-level functions if needed.
{listing}
[!htbp] {minted} [fontsize=, frame=none]cpp std::vector<double> v = 1.0, 2.0, 3.0, 4.0; Plaintext p = cc->MakeCKKSPackedPlaintext(v); auto c = cc->Encrypt(public_key, p); auto c_squared = cc->EvalMult(c, c); Plaintext result; cc->Decrypt(private_key, c_squared, &result); result->SetLength(batch_size); std::cout << "v * v = " << result; // Output: v * v = (1.0, 4.0, 9.0, … {minted} [fontsize=, frame=none]julia v = [1.0, 2.0, 3.0, 4.0] p = MakeCKKSPackedPlaintext(cc, v) c = Encrypt(cc, public_key, p) c_squared = EvalMult(cc, c, c) result = Plaintext() Decrypt(cc, private_key, c_squared, result) SetLength(result, batch_size) println("v * v = ", result) # Output: v * v = (1.0, 4.0, 9.0, … Ciphertext multiplication using OpenFHE in C++ (left) and OpenFHE.jl in Julia (right).
{listing}
[!htbp] {minted}[frame=none]julia v = [1.0, 2.0, 3.0, 4.0] c = encrypt(v, public_key, context) c_squared = c * c result = decrypt(c_squared, private_key) println("v * v = ", result) # Output: v * v = [1.0, 4.0, 9.0, … Ciphertext multiplication using SecureArithmetic.jl in Julia.
@@ 3 Accuracy and performance of the CKKS scheme
To provide a baseline for the evaluation of more complex numerical simulations using FHE, in this section we will analyze the accuracy and performance of basic CKKS operations as provided by OpenFHE. In the following, we will first describe the CKKS configuration used throughout this paper and the method we used to measure errors and runtime. We will then present our accuracy and performance analysis for individual FHE operations and discuss the implications of these results for the secure numerical simulations shown later. All code used for the experiments and our numerical results are available in the reproducibility repository [30].
@@ 3.1 Experimental setup and measurement methodology
The experiments in this paper were conducted with the OpenFHE library v1.2.0 via the SecureArithmetic.jl package, using a CKKS configuration that provides a 128-bit security level and uses $S_{\text{word}}$ = 64 bits as the native integer size. The security level is determined based on the homomorphic encryption standard [49], and OpenFHE throws an error if any parameters provided violate the security guarantee. All other parameters were chosen on a best-effort basis to balance performance and accuracy. Unless noted otherwise, the configuration used for the OpenFHE setup is as shown in Table 1. While the OpenFHE library has many more user-configurable options, we only focus on those that need to be set explicitly or where we deviate from the default values. For details on the CKKS configuration parameters, please refer to the OpenFHE documentation [40].
@@
For the batch size/capacity, we used the smallest power of two that still fits the data. The multiplicative depth available after bootstrapping was chosen to be $l_{\text{refresh}}=15$, with a maximum multiplicative depth of $l_{\text{max}}=33$. The ring dimension is chosen automatically by OpenFHE, which usually corresponds to $N_{R}=2^{17}$ given the remaining parameters. Note that the security standards in the OpenFHE library are being constantly improved, thus in the future a different ring dimension may be supported for these parameters. Furthermore, since the current homomorphic encryption standard [49] does not provide CKKS parameters when using a sparse ternary distribution for the secret key generation, in production settings it is advisable to use the uniform ternary distribution.
@@
with $L=64$. Scalars are always initialized to $301+\pi/$ 30, and plaintexts to
@@
also with a length of $L=64$. If required, a second ciphertext was created by encrypting the above plaintext. Therefore, the batch size/capacity is always $64$.
Initial tests confirmed that the computational overhead of using OpenFHE through OpenFHE.jl or SecureArithmetic.jl is negligible (below 1%). Therefore, we always used SecureArithmetic.jl to analyze the performance. The experiments were conducted on an AMD Ryzen Threadripper 3990X 64 core processor with 256 GiB RAM using a single thread. All measurements discard the initial just-in-time compilation time of the Julia programming language. We measured the runtime of FHE operations by averaging over five consecutive executions.
Numerical errors are measured by comparing the decrypted result of a secure operation with a corresponding unencrypted operation. In case of encryption, decryption, and bootstrapping operations, the error is calculated relative to doing nothing with the unencrypted data. Secure addition, multiplication, and rotation are compared to regular addition, multiplication, and Julia’s Base.circshift. For the error measurements, we used the $L^{\infty}$ norm, which denotes the maximum absolute element-wise error. Please note that the error measurements should not be interpreted by their exact numbers, but rather by their general size and trend. Different input data or different CKKS configurations may lead to different error values.
@@ 3.2 Plaintext encoding and decoding
First, we will consider the accuracy and performance of the encoding and decoding operations. As described in Sec. 2.1.1, the encoding operation is used to transform a vector of real numbers into a plaintext, where it is stored as a polynomial. The decoding is the inverse operation, which converts the polynomial back into a data vector. Unfortunately, the OpenFHE library does not provide a direct way to decode a plaintext (as extra noise is added during decoding for security purposes), thus we can only measure the runtime of the encoding operation, but not the decoding operation, nor the error of either operation. From Fig. 3, we can see that runtime required for encoding a data vector into a plaintext grows linearly with the multiplicative depth $l_{\text{max}}$. While generally the plaintext encoding is independent of $l_{\text{max}}$, by default the OpenFHE library already prepares the plaintext for multiplication with a ciphertext, i.e., transforms the plaintext to ciphertext space. This requires additional expensive operations, which depend on $l_{\text{max}}$, to be performed during encoding. Note that the level used to prepare the plaintext can be configured (changed from its default value) by the user of the library.

### v3

@@ 2.1.4 Addition
This operation results in an error growth from $e$ to $e+\hat{e}$ in decrypted data, which is still small compared to the data itself [19]. To add a plaintext to a ciphertext, the plaintext is encoded and then extended to the ciphertext space (no additional error is introduced by this operation). Addition of ciphertexts and scalars is done similarly (but no encoding is needed in this case).
@@ 2.1.5 Key Switching
Unlike addition, the two other basic arithmetic operations, multiplication and rotation, require executing a special maintenance procedure called key switching. This key switching procedure needs an additional public key, which is often called an evaluation key or key-switching hint/key. We use the term evaluation key in the paper.
@@
In CKKS, the key switching procedure is more computationally expensive than the actual multiplication or rotation. This is because key switching involves multiple number-theoretic transforms (a specialized version of the discrete Fourier transform), which dominate the overall cost of these encrypted operations in practice [28, 41]. As such, key switching is typically the main performance bottleneck in CKKS-based FHE applications.
@@ 2.1.6 Multiplication
Multiplication of two CKKS ciphertexts $\bm{c}$ and $\hat{\bm{c}}$ is a considerably more complex operation than addition, which is why we only provide a brief overview here. It requires three steps:
@@
As we saw in Sec. 2.1.3, after encryption the first polynomial of the ciphertext $c_{0}$ is linear with respect to the secret key $\bm{\mathrm{sk}}$. After multiplying two ciphertexts using the tensor product, we get a quadratic polynomial in secret key polynomial $s$, i.e., the decryption would be evaluated as $\langle\bm{d},(1,s,s^{2})\rangle=d_{0}+d_{1}\cdot s+d_{2}\cdot s^{2}$. Each subsequent multiplication will increase the degree of $s$ even further; one can think of a multiplication of two ciphertexts as a product of two corresponding decryption polynomials. This implies the size of the ciphertext after a multiplication of two ciphertexts of sizes $i$ and $j$ will grow to i+j1i+j-1 polynomials. To maintain the compact representation of ciphertexts, the relinearization procedure compresses the ciphertext by reducing it by one polynomial, in this case from three to two polynomials.
@@
In leveled FHE, one therefore tries to minimize the required multiplicative depth of an algorithm by evaluating a chain of multiplications using a binary tree approach. For example, instead of computing $\bm{c^{\prime}}=\bm{c}_{1}\cdot\bm{c}_{2}\cdot\bm{c}_{3}\cdot\bm{c}_{4}$ with three consecutive multiplications (and thus requiring three levels), one can first compute $\bm{r}_{1}=\bm{c}_{1}\cdot\bm{c}_{2}$ and $\bm{r}_{2}=\bm{c}_{3}\cdot\bm{c}_{4}$ using one level, and then combine them to $\bm{c^{\prime}}=\bm{r}_{1}\cdot\bm{r}_{2}$ using a second level. In the naïve case, to multiply $n$ ciphertexts, the multiplicative depth of n1n-1 is needed. If the binary tree method is used instead, the multiplicative depth is reduced to approximately $\log_{2}(n)$ (see also [38]).
@@ 2.1.7 Rotation
The multiplication and addition operations are element-wise over the entire ciphertext, i.e., they are always performed between corresponding elements of the user-provided data vectors. However, many practical algorithms require interactions between elements within the same vector or access to specific indices, which goes beyond basic element-wise operations. In such cases, the rotation operation is used, which cyclically shifts encrypted data by some rotation index. When the rotation is evaluated, the underlying secret key is also implicitly rotated. To change the underlying secret key back to the original one (to support further homomorphic computations), one needs to apply the key switching operation. The rotation itself is cheap (effectively it is just reindexing), but the required key switching operation following it is computationally expensive.
@@ 2.1.8 Bootstrapping
As described in Sec. 2.1.2 and 2.1.6, the necessity to rescale the ciphertext limits the number of multiplicative operations that can be performed, since no more multiplications are possible once the level $l$ reaches zero. The bootstrapping operation alleviates this restriction by resetting the level to $l>0$. It involves approximating a homomorphic decryption procedure on a ciphertext using an encrypted version of the secret key (this is done indirectly in CKKS), and then re-encrypting it to obtain a refreshed ciphertext. In theory, the number of multiplications that can be performed with a ciphertext is unlimited when using bootstrapping. However, bootstrapping also adds a significant approximation error to the ciphertext and is computationally very expensive. Due to the high complexity of the bootstrapping operation, we do not go into more detail here. In-depth descriptions of various bootstrapping implementations are covered elsewhere [42].
@@ 2.2 Software libraries for fully homomorphic encryption
While the CKKS scheme is a powerful method for FHE with real numbers, it is also an advanced technique that requires a deep understanding of the underlying mathematical methods to be used effectively. It is therefore advisable to use a software library that provides an implementation of the CKKS scheme and other necessary operations. Several such libraries are available, e.g., SEAL [43], HElib [44], and OpenFHE [40].
In this paper, we focus on the OpenFHE library, which is actively maintained and includes most recent CKKS optimizations [45]. We also present two Julia packages OpenFHE.jl and SecureArithmetic.jl, which provide a convenient interface to OpenFHE in the Julia programming language. We now give a brief overview of these tools.
@@ 2.2.1 OpenFHE
OpenFHE [40] is an open-source FHE library that offers efficient C++ implementations of all common FHE schemes, including the CKKS scheme with bootstrapping. OpenFHE contains many examples that are especially useful for new users when designing their applications. The OpenFHE library puts an emphasis on usability. For example, the relinearization and rescaling operations after ciphertext multiplication are handled automatically, allowing the library to be used also by non-experts in FHE. This can be observed in a short code snippet in Listing 3 (left), where the CKKS internals for multiplication are hidden from the user. Some code for initializing the setup was omitted for clarity, e.g., to create the cryptographic context object (named cc in this listing) or the public and private keys. For more details about OpenFHE, please refer to [40].
@@
It is this feature that we leverage in our OpenFHE.jl package [47], which provides a Julia wrapper for the OpenFHE library. When installing OpenFHE.jl, pre-built binaries of the OpenFHE library are automatically downloaded without the user having to compile anything locally. The C++ functionality is exposed in Julia via the CxxWrap.jl package [48]. Besides offering a native Julia interface to the OpenFHE library, OpenFHE.jl does not provide any extra functionality. It is intended to be used as a backend for other Julia packages that require FHE capabilities. As can be seen in Listing 3 (right), the syntax of OpenFHE.jl is very similar to that of OpenFHE, making it easy to port code between the two programming languages.
@@
To make the use of FHE more accessible to non-experts in Julia, we created the SecureArithmetic.jl package [49]. It inherits the low-level functionalities of the OpenFHE.jl package, but provides them through a more convenient, higher-level Julia interface. Using the SecureArithmetic.jl package, users can write arithmetic expressions in a common mathematical notation, significantly simplifying the code. Furthermore, SecureArithmetic.jl supports the ability to use the same code for secure and non-secure computations, which is especially useful during the design stage. It allows one to debug FHE algorithms without having to go through the actual encryption and decryption steps, making the execution much faster during the prototyping phase. Listing 4 shows an example of how the ciphertext multiplication is performed using SecureArithmetic.jl. Compared to OpenFHE/OpenFHE.jl in Listing 3, SecureArithmetic.jl provides a more high-level interface that is closer to the mathematical notation. However, since it is built on top of OpenFHE.jl and OpenFHE, it is still possible to access the lower-level functions if needed.
std::vector<double> v = {1.0, 2.0, 3.0, 4.0};
Plaintext p = cc->MakeCKKSPackedPlaintext(v);
auto c = cc->Encrypt(public_key, p);
auto c_squared = cc->EvalMult(c, c);
Plaintext result;
cc->Decrypt(private_key, c_squared, &result);
result->SetLength(batch_size);
std::cout << "v * v = " << result;
// Output: v * v = (1.0, 4.0, 9.0,
v = [1.0, 2.0, 3.0, 4.0]
p = MakeCKKSPackedPlaintext($c$, v)
c = Encrypt($c$, public_key, p)
c_squared = EvalMult($c$, c, c)
result = Plaintext()
Decrypt($c$, private_key, c_squared, result)
SetLength(result, batch_size)
println("v * v = ", result)
# Output: v * v = (1.0, 4.0, 9.0,
Figure 3: Ciphertext multiplication using OpenFHE in C++ (left) and OpenFHE.jl in Julia (right).
v = [1.0, 2.0, 3.0, 4.0]
c = encrypt(v, public_key, context)
c_squared = c * c
result = decrypt(c_squared, private_key)
println("v * v = ", result)
# Output: v * v = [1.0, 4.0, 9.0,
Figure 4: Ciphertext multiplication using SecureArithmetic.jl in Julia.
@@ 3 Accuracy and performance of the CKKS scheme
In this section, we analyze the accuracy and performance of basic CKKS operations as provided by OpenFHE to create a baseline for the more complex numerical simulations shown later. We first describe the CKKS configuration used throughout this paper and the method we use to measure errors and runtime. We then present our accuracy and performance analysis for individual FHE operations and discuss the implications of these results for the secure numerical simulations. All code used for the experiments and our numerical results are available in the reproducibility repository [30].
@@ 3.1 Experimental setup and measurement methodology
The experiments in this paper were conducted with the OpenFHE library v1.2.0 via the SecureArithmetic.jl package, using a CKKS configuration that ensures 128-bit security and uses $S_{\text{word}}$ = 64 bits as the native integer size. The security level is determined based on the homomorphic encryption standard [50], and OpenFHE throws an error if any parameters provided violate the security guarantee. All other parameters were chosen on a best-effort basis to balance performance and accuracy. Unless noted otherwise, the configuration used for the OpenFHE setup is as shown in Table 1. While the OpenFHE library has many more user-configurable options, we only focus on those that need to be set explicitly or where we deviate from the default values. For details on the CKKS configuration parameters, please refer to the OpenFHE documentation [40].
@@
For the batch size/capacity, we used the smallest power of two that still fits the data. The multiplicative depth available after bootstrapping was chosen to be $l_{\text{refresh}}=15$, with a maximum multiplicative depth of $l_{\text{max}}=33$. The ring dimension is chosen automatically by OpenFHE, which usually corresponds to $N_{R}=2^{17}$ given the remaining parameters. Since the security standards in the OpenFHE library are being constantly improved, a different ring dimension may be supported for these parameters in the future. Furthermore, since the current homomorphic encryption standard [50] does not provide CKKS parameters when using a sparse ternary distribution for the secret key generation, in production settings it is advisable to use the uniform ternary distribution.
@@
with $L=64$. Scalars are always initialized to $1+\pi/30$, and plaintexts to
@@
also with a length of $L=64$. Therefore, the batch size/capacity is always $64$. If required, a second ciphertext was created by encrypting the above plaintext. Due to the approximate nature of CKKS, the encrypted values differ slightly from the original.
Since the computational overhead of using OpenFHE through the Julia packages is negligible compared to the cost of the FHE operations themselves, in this paper we use SecureArithmetic.jl to analyze the performance. The experiments were conducted on an AMD Ryzen Threadripper 3990X 64 core processor with 256 GiB RAM using a single thread. All measurements discard the initial just-in-time compilation time of the Julia programming language. We measured the runtime of FHE operations by averaging over five consecutive executions.
Numerical errors are measured by comparing the decrypted result of a secure operation with a corresponding unencrypted operation. In case of encryption, decryption, and bootstrapping operations, the error is calculated relative to doing nothing with the unencrypted data. Secure addition, multiplication, and rotation are compared to regular addition, multiplication, and Julia’s Base.circshift. For the error measurements, we used the $L^{\infty}$ norm, which denotes the maximum absolute element-wise error. The observed error depends heavily on the cryptographic parameters such as ring dimension, security level, or scaling modulus. Therefore, error measurements should not be interpreted by their absolute numbers, but rather by their order of magnitude and trend.
@@ 3.2 Plaintext encoding and decoding
As described in Sec. 2.1.1, the encoding operation is used to transform a vector of real numbers into a plaintext, where it is stored as a polynomial. The decoding is the inverse operation, which converts the polynomial back into a data vector. Unfortunately, the OpenFHE library does not provide a direct way to decode a plaintext (as extra noise is added during decoding for security purposes), thus we can only measure the runtime of the encoding operation, but not the decoding operation, nor the error of either operation. From Fig. 5, we can see that runtime required for encoding a data vector into a plaintext grows linearly with the multiplicative depth $l_{\text{max}}$. While generally the plaintext encoding is independent of $l_{\text{max}}$, by default the OpenFHE library already prepares the plaintext for multiplication with a ciphertext, i.e., transforms the plaintext to ciphertext space. This requires additional expensive operations, which depend on $l_{\text{max}}$, to be performed during encoding. The level used to prepare the plaintext can be configured (changed from its default value) by the user of the library.